# Where does the simplified Stellar Contamination Model fail in Exoplanet Transmission Spectroscopy?

### v1

@@ 3.1 Idealised model–model comparison
Importantly, recent JWST eclipse observations of WASP-69 b also support the presence of high-altitude aerosols on the dayside hemisphere (Schlawin.et.al.2024AJ....168..104S). Although those emission spectra are unaffected by stellar heterogeneity, the fact that cloud-free models fail to reproduce the observed flux reinforces the interpretation that aerosols, rather than stellar contamination, likely shape the near-infrared features. This highlights the need for transmission analyses to account for both stellar and planetary contributions explicitly, as done in ECLIPSE-X, which treats stellar contamination independently of atmospheric retrievals but allows for their disentanglement across wavelengths.
@@ 3.2 Application to observational data: LHS 1140 b
The transmission spectrum analysed in this section is the combined JWST/NIRISS SOSS data set of LHS 1140 b, consisting of two transits reduced and binned to (Cadieux.et.al.2024ApJ...970L...2C)333The observational spectrum was provided by the authors of Cadieux.et.al.2024ApJ...970L...2C. For reference, the R–TLSE column in Table 2 reproduces the TLS-only stellar-contamination solution reported by Cadieux.et.al.2024ApJ...970L...2C at this resolution. We then fit the same spectrum with ECLIPSE-X by exploring a regular grid in the temperatures and filling factors of spots and faculae. For each grid point, ECLIPSE-X computes the corresponding disc-integrated photosphere + activity spectrum, folds it through the NIRISS/SOSS throughput and spectral response to match the channel grid, incorporates the transit light source (TLS) effect in the model transit depths, and evaluates the resulting transmission spectrum against the observations via a statistic. The best-fitting solutions for models with our nominal limb-darkening treatment and with limb darkening set to zero (LDCs = 0) are shown in Figure 5 and summarised in Table 2.
@@
As a first step, we compare the ECLIPSE-X retrieval with limb darkening switched off (LDCs = 0) to the reference R–TLSE solution. When all limb-darkening coefficients are set to zero, the stellar disc becomes uniformly bright and the transit chord no longer samples any centre-to-limb intensity gradient. In this case, the pixelised ECLIPSE-X simulations approach the same disc-averaged description of stellar contamination that underlies the R–TLSE formalism. This modelling configuration is closely related to the situation discussed by Alexoudi.et.al.2020A&A...640A.134A, who showed that removing the host-star limb darkening causes the impact-parameter degeneracy to disappear: the inferred planet-to-star radius ratio no longer exhibits any wavelength-dependent offset and remains constant across the spectrum. Adopting this limb-darkening-free configuration in ECLIPSE-X therefore eliminates that particular source of bias and isolates any residual differences between our model and R–TLSE.
@@
This behaviour is readily understood once we account for the impact of limb darkening on the baseline transit geometry. Even in the absence of stellar heterogeneities, a limb-darkened star does not, in general, produce a perfectly flat transmission spectrum: for a nearly central transit, the combination of impact parameter and centre-to-limb intensity variations naturally induces a modest positive slope in at optical wavelengths, as highlighted by Alexoudi.et.al.2020A&A...640A.134A. In the R–TLSE (or LDCs = 0) limit, this geometric effect is explicitly removed by construction, because the stellar disc is assumed to be uniformly bright, and the baseline spectrum is flat in the absence of active regions. Unocculted faculae then act on top of this flat reference and directly imprint a net slope. In the fully limb-darkened ECLIPSE-X framework, on the other hand, we already start from a positively sloped baseline set by the transit geometry; adding unocculted faculae must first compensate for this geometrically induced trend before it can reproduce any additional slope of opposite sign. The optimiser therefore tends to increase and to push towards the upper bound of the prior in order to counteract the limb-darkening-driven slope. However, foreshortening and the localised nature of the active regions limit how much extra bias they can introduce into the disc-integrated flux. As a result, R–TLSE-like slopes can be reproduced in the fully limb-darkened simulations only in the extreme regime of very hot faculae covering –40% of the visible hemisphere. More moderate facular configurations naturally produce a shallower contamination signal, implying that any residual optical slope is likely to include a genuine atmospheric contribution, potentially combined with a less extreme level of stellar contamination, rather than being explained solely by unocculted faculae.
@@ 4 Summary and Conclusions
This study demonstrates that the Rackham.et.al.2018ApJ...853..122R Transit Light Source Effect (R–TLSE) approximation introduces systematic, wavelength-dependent biases in exoplanet transmission spectra, especially at optical wavelengths where limb-darkening gradients and transit geometry play a dominant role. Through idealised, noise-free simulations (Section 3.1), we compare the disc-averaged R–TLSE correction to our self-consistent ECLIPSE-X framework, which incorporates wavelength-dependent limb darkening, the spatial distribution of active regions, and the full transit geometry, and we quantify system-specific discrepancies for three archetypal planets. We then apply both approaches to the JWST/NIRISS SOSS observations of LHS 1140 b (Section 3.2), using the data as a testbed for assessing the practical impact of these differences on atmospheric interpretation. Overall, our analysis shows that the scale of the R–TLSE-induced errors must be assessed on a case-by-case basis.
For LHS 1140 b, a super-Earth orbiting an M dwarf, we find peak discrepancies of 170 ppm due to starspots and 140 ppm due to faculae at 500 nm, driven primarily by the R–TLSE’s neglect of optical limb-darkening contrasts. These errors decrease to below 10 ppm at near-infrared (NIR) wavelengths above 1000 nm, where limb-darkening effects are weaker, rendering R–TLSE-based retrievals more reliable. However, in the optical regime (500–800 nm), R–TLSE inaccuracies become largest and may contribute to marginal atmospheric detections. Notably, the spectral region supporting the tentative N2 signal reported by Cadieux.et.al.2024ApJ...970L...2C coincides with wavelengths where R–TLSE inaccuracies are expected to peak and with the part of the spectrum that exhibits the largest observational uncertainties and dispersion, making any N2 claim especially sensitive to the adopted stellar-contamination model.
@@ Acknowledgements.
VYDS and AV acknowledge the partial financial support received from Brazilian FAPESP grants #2021/14897-9, #2018/04055-8 #2021/02120-0, and #2024/03652-3, as well as CAPES and MackPesquisa funding agencies. This research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). We thank the authors of Cadieux.et.al.2024ApJ...970L...2C for kindly providing the observational data of LHS 1140 b used in this work.Appendix A Supplementary Figures

### v2

@@ 3.1 Idealised model–model comparison
Importantly, recent JWST eclipse observations of WASP-69 b also support the presence of high-altitude aerosols on the dayside hemisphere (Schlawin et al., 2024). Although those emission spectra are unaffected by stellar heterogeneity, the fact that cloud-free models fail to reproduce the observed flux reinforces the interpretation that aerosols, rather than stellar contamination, likely shape the near-infrared features. This highlights the need for transmission analyses to account for both stellar and planetary contributions explicitly, as done in ECLIPSE-X, which treats stellar contamination independently of atmospheric retrievals but allows for their disentanglement across wavelengths.
@@ 3.2 Application to observational data: LHS 1140 b
The transmission spectrum analysed in this section is the combined JWST/NIRISS SOSS data set of LHS 1140 b, consisting of two transits reduced and binned to (Cadieux et al., 2024)333The observational spectrum was provided by the authors of Cadieux et al. (2024). For reference, the R–TLSE column in Table 2 reproduces the TLS-only stellar-contamination solution reported by Cadieux et al. (2024) at this resolution. We then fit the same spectrum with ECLIPSE-X by exploring a regular grid in the temperatures and filling factors of spots and faculae. For each grid point, ECLIPSE-X computes the corresponding disc-integrated photosphere + activity spectrum, folds it through the NIRISS/SOSS throughput and spectral response to match the channel grid, incorporates the transit light source (TLS) effect in the model transit depths, and evaluates the resulting transmission spectrum against the observations via a statistic. The best-fitting solutions for models with our nominal limb-darkening treatment and with limb darkening set to zero (LDCs = 0) are shown in Figure 5 and summarised in Table 2.
@@
As a first step, we compare the ECLIPSE-X retrieval with limb darkening switched off (LDCs = 0) to the reference R–TLSE solution. When all limb-darkening coefficients are set to zero, the stellar disc becomes uniformly bright and the transit chord no longer samples any centre-to-limb intensity gradient. In this case, the pixelised ECLIPSE-X simulations approach the same disc-averaged description of stellar contamination that underlies the R–TLSE formalism. This modelling configuration is closely related to the situation discussed by Alexoudi et al. (2020), who showed that removing the host-star limb darkening causes the impact-parameter degeneracy to disappear: the inferred planet-to-star radius ratio no longer exhibits any wavelength-dependent offset and remains constant across the spectrum. Adopting this limb-darkening-free configuration in ECLIPSE-X therefore eliminates that particular source of bias and isolates any residual differences between our model and R–TLSE.
@@
This behaviour is readily understood once we account for the impact of limb darkening on the baseline transit geometry. Even in the absence of stellar heterogeneities, a limb-darkened star does not, in general, produce a perfectly flat transmission spectrum: for a nearly central transit, the combination of impact parameter and centre-to-limb intensity variations naturally induces a modest positive slope in at optical wavelengths, as highlighted by Alexoudi et al. (2020). In the R–TLSE (or LDCs = 0) limit, this geometric effect is explicitly removed by construction, because the stellar disc is assumed to be uniformly bright, and the baseline spectrum is flat in the absence of active regions. Unocculted faculae then act on top of this flat reference and directly imprint a net slope. In the fully limb-darkened ECLIPSE-X framework, on the other hand, we already start from a positively sloped baseline set by the transit geometry; adding unocculted faculae must first compensate for this geometrically induced trend before it can reproduce any additional slope of opposite sign. The optimiser therefore tends to increase and to push towards the upper bound of the prior in order to counteract the limb-darkening-driven slope. However, foreshortening and the localised nature of the active regions limit how much extra bias they can introduce into the disc-integrated flux. As a result, R–TLSE-like slopes can be reproduced in the fully limb-darkened simulations only in the extreme regime of very hot faculae covering –40% of the visible hemisphere. More moderate facular configurations naturally produce a shallower contamination signal, implying that any residual optical slope is likely to include a genuine atmospheric contribution, potentially combined with a less extreme level of stellar contamination, rather than being explained solely by unocculted faculae.
@@ 4 Summary and conclusions
This study demonstrates that the Rackham et al. (2018) Transit Light Source Effect (R–TLSE) approximation introduces systematic, wavelength-dependent biases in exoplanet transmission spectra, especially at optical wavelengths where limb-darkening gradients and transit geometry play a dominant role. Through idealised, noise-free simulations (Section 3.1), we compare the disc-averaged R–TLSE correction to our self-consistent ECLIPSE-X framework, which incorporates wavelength-dependent limb darkening, the spatial distribution of active regions, and the full transit geometry, and we quantify system-specific discrepancies for three archetypal planets. We then apply both approaches to the JWST/NIRISS SOSS observations of LHS 1140 b (Section 3.2), using the data as a testbed for assessing the practical impact of these differences on atmospheric interpretation. Overall, our analysis shows that the scale of the R–TLSE-induced errors must be assessed on a case-by-case basis.
For LHS 1140 b, a super-Earth orbiting an M dwarf, we find peak discrepancies of 170 ppm due to starspots and 140 ppm due to faculae at 500 nm, driven primarily by the R–TLSE’s neglect of optical limb-darkening contrasts. These errors decrease to below 10 ppm at near-infrared (NIR) wavelengths above 1000 nm, where limb-darkening effects are weaker, rendering R–TLSE-based retrievals more reliable. However, in the optical regime (500–800 nm), R–TLSE inaccuracies become largest and may contribute to marginal atmospheric detections. Notably, the spectral region supporting the tentative N2 signal reported by Cadieux et al. (2024) coincides with wavelengths where R–TLSE inaccuracies are expected to peak and with the part of the spectrum that exhibits the largest observational uncertainties and dispersion, making any N2 claim especially sensitive to the adopted stellar-contamination model.
@@ Acknowledgements.
VYDS and AV acknowledge the partial financial support received from Brazilian FAPESP grants #2021/14897-9, #2018/04055-8 #2021/02120-0, and #2024/03652-3, as well as CAPES and MackPesquisa funding agencies. This research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). We thank the authors of Cadieux et al. (2024) for kindly providing the observational data of LHS 1140 b used in this work.References
- Role of the impact parameter in exoplanet transmission spectroscopy. A&A 640, pp. A134. External Links: Document, 2007.07716, ADS entry Cited by: §3.1, §3.1, §3.2, §3.2.
- Kepler-411 Differential Rotation from Three Transiting Planets. ApJ 907 (1), pp. L5. External Links: Document, 2101.07692, ADS entry Cited by: §2.1.
- Hubble Space Telescope Time-Series Photometry of the Transiting Planet of HD 209458. ApJ 552 (2), pp. 699–709. External Links: Document, astro-ph/0101336, ADS entry Cited by: §1.
- Transmission Spectroscopy of the Habitable Zone Exoplanet LHS 1140 b with JWST/NIRISS. ApJ 970 (1), pp. L2. External Links: Document, 2406.15136, ADS entry Cited by: §1, Figure 5, Figure 5, 1st item, 2nd item, §3.1, §3.1, §3.2, §4, §4, footnote 3, footnote 4.
- A Hubble PanCET Study of HAT-P-11b: A Cloudy Neptune with a Low Atmospheric Metallicity. AJ 158 (6), pp. 244. External Links: Document, 1910.07523, ADS entry Cited by: §1.
- Detection of an Extrasolar Planet Atmosphere. ApJ 568 (1), pp. 377–384. External Links: Document, astro-ph/0111544, ADS entry Cited by: §1.
- Limb-darkening coefficients for four-term and power-2 laws for the JWST mission adopting spherical PHOENIX models at high resolution: NIRCam, NIRISS, and NIRSpec passbands. A&A 699, pp. A97. External Links: Document, 2506.07265, ADS entry Cited by: §2.2.
- A new non-linear limb-darkening law for LTE stellar atmosphere models. Calculations for -5.0 ¡= log[M/H] ¡= +1, 2000 K ¡= Teff ¡= 50000 K at several surface gravities. A&A 363, pp. 1081–1190. External Links: ADS entry Cited by: §2.2.
- Detection of Aerosols at Microbar Pressures in an Exoplanet Atmosphere. AJ 162 (3), pp. 91. External Links: Document, 2106.10292, ADS entry Cited by: §3.1.
- Near-infrared transmission spectroscopy of HAT-P-18 b with NIRISS: Disentangling planetary and stellar features in the era of JWST. MNRAS 528 (2), pp. 3354–3377. External Links: Document, 2310.14950, ADS entry Cited by: §1.
- Numerical solution of the expanding stellar atmosphere problem.. Journal of Computational and Applied Mathematics 109 (1), pp. 41–63. External Links: Document, astro-ph/9808182, ADS entry Cited by: §2.2.
- Parallel Implementation of the PHOENIX Generalized Stellar Atmosphere Program. ApJ 483 (1), pp. 390–398. External Links: Document, astro-ph/9607087, ADS entry Cited by: §2.2.
- Modelling the photosphere of active stars for planet detection and characterization. A&A 586, pp. A131. External Links: Document, 1511.06717, ADS entry Cited by: §1.
- A new extensive library of PHOENIX stellar atmospheres and synthetic spectra. A&A 553, pp. A6. External Links: Document, 1303.5632, ADS entry Cited by: §2.2.
- The Influence of Stellar Contamination on the Interpretation of Near-infrared Transmission Spectra of Sub-Neptune Worlds around M-dwarfs. ApJ 889 (2), pp. 78. External Links: Document, 1912.04389, ADS entry Cited by: §1, §3.1.
- Atmospheric Reconnaissance of TRAPPIST-1 b with JWST/NIRISS: Evidence for Strong Stellar Contamination in the Transmission Spectra. ApJ 955 (1), pp. L22. External Links: Document, 2309.07047, ADS entry Cited by: §1.
- Carbon-bearing Molecules in a Possible Hycean Atmosphere. ApJ 956 (1), pp. L13. External Links: Document, 2309.05566, ADS entry Cited by: §3.1, §3.1.
- Double Trouble: Two Transits of the Super-Earth GJ 1132 b Observed with JWST NIRSpec G395H. ApJ 959 (1), pp. L9. External Links: Document, 2310.10711, ADS entry Cited by: §1, 1st item, 2nd item.
- Water Vapor in the Spectrum of the Extrasolar Planet HD 189733b. I. The Transit. ApJ 791 (1), pp. 55. External Links: Document, 1407.2462, ADS entry Cited by: §1.
- Using the Sun to estimate Earth-like planets detection capabilities . II. Impact of plages. A&A 512, pp. A39. External Links: Document, 1001.1638, ADS entry Cited by: 2nd item.
- High Tide or Riptide on the Cosmic Shoreline? A Water-rich Atmosphere or Stellar Contamination for the Warm Super-Earth GJ 486b from JWST Observations. ApJ 948 (1), pp. L11. External Links: Document, 2305.00868, ADS entry Cited by: §1.
- The GTC exoplanet transit spectroscopy survey. XI. Possible detection of Rayleigh scattering in the atmosphere of the Saturn-mass planet WASP-69b. A&A 641, pp. A158. External Links: Document, 2007.02741, ADS entry Cited by: §3.1.
- Stellar magnetic activity and the butterfly diagram of Kepler-63. A&A 635, pp. A78. External Links: Document, 1911.08661, ADS entry Cited by: §2.1.
- LDTK: Limb Darkening Toolkit. MNRAS 453 (4), pp. 3821–3826. External Links: Document, 1508.02634, ADS entry Cited by: Figure 6, Figure 6, §2.2.
- The Influence of Stellar Chromospheres and Coronae on Exoplanet Transmission Spectroscopy. ApJ 980 (2), pp. L42. External Links: Document, 2502.00553, ADS entry Cited by: §1.
- Detection of faculae in the transit and transmission spectrum of WASP-69b. A&A 692, pp. A83. External Links: Document, 2410.18663, ADS entry Cited by: §1, §3.1, §3.1.
- ACCESS I: An Optical Transmission Spectrum of GJ 1214b Reveals a Heterogeneous Stellar Photosphere. ApJ 834 (2), pp. 151. External Links: Document, 1612.00228, ADS entry Cited by: §1.
- The Transit Light Source Effect: False Spectral Features and Incorrect Densities for M-dwarf Transiting Planets. ApJ 853 (2), pp. 122. External Links: Document, 1711.05691, ADS entry Cited by: §1, §1, §1, §2.3, §2.3, §4.
- The Transit Light Source Effect. II. The Impact of Stellar Heterogeneity on Transmission Spectra of Planets Orbiting Broadly Sun-like Stars. AJ 157 (3), pp. 96. External Links: Document, 1812.06184, ADS entry Cited by: §1, §2.3.
- Toward Robust Corrections for Stellar Contamination in JWST Exoplanet Transmission Spectra. AJ 168 (2), pp. 82. External Links: Document, 2303.15418, ADS entry Cited by: §3.1.
- Promise and Peril: Stellar Contamination and Strict Limits on the Atmosphere Composition of TRAPPIST-1 c from JWST NIRISS Transmission Spectra. ApJ 979 (1), pp. L5. External Links: Document, 2409.19333, ADS entry Cited by: §1.
- A Population Analysis of 20 Exoplanets Observed from Optical to Near-infrared Wavelengths with the Hubble Space Telescope: Evidence for Widespread Stellar Contamination. ApJS 276 (2), pp. 70. External Links: Document, 2404.15505, ADS entry Cited by: §1.
- Multiple Clues for Dayside Aerosols and Temperature Gradients in WASP-69 b from a Panchromatic JWST Emission Spectrum. AJ 168 (3), pp. 104. External Links: Document, 2406.15543, ADS entry Cited by: §3.1.
- A Comprehensive Reanalysis of K2-18 b’s JWST NIRISS+NIRSpec Transmission Spectrum. AJ 170 (6), pp. 298. External Links: Document, 2501.18477, ADS entry Cited by: §3.1.
- Theoretical Transmission Spectra during Extrasolar Giant Planet Transits. ApJ 537 (2), pp. 916–921. External Links: Document, astro-ph/9912241, ADS entry Cited by: §1.
- Method for Spot Detection on Solar-like Stars. ApJ 585 (2), pp. L147–L150. External Links: Document, ADS entry Cited by: §2.1, §2.1.
- Properties of starspots on CoRoT-2. A&A 510, pp. A25. External Links: Document, 0909.4055, ADS entry Cited by: §2.1.
- ECLIPSE-xlambda. Zenodo. External Links: Link, Document Cited by: §2.1.
- Differential Rotation of CoRoT Stars and a Kepler Binary Star from Starspot Transit Mapping. ApJ 972 (1), pp. 81. External Links: Document, ADS entry Cited by: §2.1.
- Stellar Obliquity from Spot Transit Mapping of Kepler-210. ApJ 940 (2), pp. 132. External Links: Document, ADS entry Cited by: §2.1.
- Activity and Rotation of Kepler-17. ApJ 835 (2), pp. 294. External Links: Document, 1702.02213, ADS entry Cited by: §2.1.
- Activity and differential rotation of the early M dwarf Kepler-45 from transit mapping. MNRAS 492 (4), pp. 5141–5151. External Links: Document, ADS entry Cited by: §2.1.
- Differential rotation of Kepler-71 via transit photometry mapping of faculae and starspots. MNRAS 484 (1), pp. 618–630. External Links: Document, ADS entry Cited by: §2.1.
- Dynamo activity of the k dwarf koi-883 from transit photometry mapping. Monthly Notices of the Royal Astronomical Society 510 (4), pp. 5348–5361. Cited by: §2.1.
- The Near-infrared Transmission Spectra of TRAPPIST-1 Planets b, c, d, e, f, and g and Stellar Contamination in Multi-epoch Transit Spectra. AJ 156 (4), pp. 178. External Links: Document, 1802.02086, ADS entry Cited by: §3.1.
Appendix A Supplementary Figures